OpenStudy (anonymous):

Identify the indeterminate form and evaluate the limit using l'Hopital's Rule. The limit as x approaches infinity of ln(x+1)/logbase2 of x

OpenStudy (psymon):

Well, if x is approaching infinity, then straight away we have theindeterminant form infinity over infinity. Since we have the function in an undeterminant form, this satisfies the condition in which we can use l'hopitals rule. Do you know the derivatives of natural log and of logs with bases other than e?

OpenStudy (anonymous):

Yeah I know the derivative of logs

OpenStudy (psymon):

Alright, so just take the derivative of the numerator and then divide it by the derivative of the denominator and let's see what we get.

OpenStudy (anonymous):

so for the bottom the derivative would be 1 / xln2 ?

OpenStudy (psymon):

Correct.

OpenStudy (anonymous):

do i do the reciprocal of the bottom derivative in order to multiply it by the top derivative ?

OpenStudy (psymon):

Right, so the derivative for the bottom would basically flip and multiply.

OpenStudy (anonymous):

so after i get \[\lim_{x \rightarrow \infty}\frac{ x \ln 2 }{ x+1 }\] what do I do ?

OpenStudy (psymon):

Well, we still have an indeterminant form. This means we can just use l'hopitals again.

OpenStudy (anonymous):

I use product rule for the top right ?

OpenStudy (psymon):

No need. ln2 is just a constant. So you take the derivative of it in the same way you would like 2x.

OpenStudy (anonymous):

so the limit would just be ln2 ?

OpenStudy (psymon):

Yep, thats all it would be :3

OpenStudy (anonymous):

Thank you :)

OpenStudy (psymon):

Sure ^_^